# Continuous-Time Quantum Markov Chains And Discretizations Of $p$-Adic Schrödinger Equations: Comparisons And Simulations

### v1

@@ 1. Introduction
Quantum computing is a computing paradigm that uses quantum physics effects like superposition, interference, and entanglement to achieve computational power beyond classical computing; see, e.g., the recent survey [1], and the references therein. The continuous-time quantum walks (CTQWs) are the quantum counterparts of the classical random walks. CTQWs are a fundamental tool for developing efficient quantum algorithms, simulating complex quantum systems, and potentially achieving quantum advantages in various computational tasks. CTQWs offer potential advantages over classical algorithms, particularly in graph-based problems like spatial search, [2]-[9]. Currently, there is a large variety of quantum physical systems that can be used for implementing quantum walks and their applications. For instance, trapped ion and trapped-atom systems, bulk optics systems, and integrated photonics systems. These considerable advances in physical implementations of quantum walks are allowing the implementation of specialized quantum walk computing systems for applications of practical interest, [1], [10].
Ultrametricity in physics means the emergence of ultrametric spaces in physical models. A metric space $(M,d)$ is ultrametric if  the distance $d$ satisfies $d(A,B)\leq\max\left\{d(A,C),d(C,B)\right\}$ for any three points $A$, $B,$ $C$ in $M$. Ultrametricity was discovered in the 80s by Parisi and others in the theory of spin glasses and by Frauenfelder and others in physics of proteins; see e.g. [11], [12], see also [13, Chapter 4], and  the references therein. In both cases, the space of states of a complex system has a hierarchical structure which play a central role the physical behavior of the system.
@@
In [26], the first author showed that a large class of $p$-adic Schrödinger equations is the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). Practically, a discretization of such an equation gives a CTQMC. As a practical result, we construct new types of continuous-time quantum walks (CTQWs) on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. The final goal of the mentioned work was to contribute in the understanding of he role of the hypothesis of the discreteness of space in the foundations of quantum mechanics. The connection between $p$-adic QM and CTQWs show that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory because the Hamiltonians used are non-local operators, and consequently, spooky actions at a distance are allowed. But, this theory is not a mathematical toy. The paradigm asserting that the universe in not locally real, implies that $p$-adic QM allows realism.
@@
$\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right),\end{array}\right$.
where $i=$ $\sqrt{-1}$. The solution of is given by $\Psi\left(x,t\right)=Z_{0}\left(x,it\right)\ast\psi_{0}\left(x\right)$, cf. Theorem 11.1, in Appendix C. Now, we construct a quantum Markov chain in $G_{l}$ using the following transition probabilities:
@@
Set $\boldsymbol{J}\varphi=J\ast\varphi-\varphi$, then the operators $\boldsymbol{J}$, $e^{t\boldsymbol{J}}$, $e^{it\boldsymbol{J}}$ are limits of finite dimensional operators $\boldsymbol{J}^{\left(l\right)}$, $e^{t\boldsymbol{J}^{\left(l\right)}}$, $e^{it\boldsymbol{J}^{\left(l\right)}}$, and consequently in the computation of $p_{r,v}(t)$, respectively of $\pi_{r,v}\left(t\right)$, we can use $Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=e^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)$, respectively $Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=e^{it\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)$, and thus the construction of the above-mentioned Markov chains depends on a matrix $\boldsymbol{J}^{\left(l\right)}$. By suitably selecting this matrix, we construct the CTQWs based on adjacency matrices of graphs, see Appendix E.
@@ 2.3. $p$-Adic QM
In the Dirac-Von Neumann formulation of QM, to every isolated quantum system there is associated a separable complex Hilbert space $\mathcal{H}$ called the space of states. The Hilbert space of a composite system is the Hilbert space tensor product of the state spaces associated with the component systems. Non-zero vectors describe the states of a quantum system from $\mathcal{H}$. Each observable corresponds to a unique linear self-adjoint operator in $\mathcal{H}$.  The most important observable of a quantum system is its energy. We denote the corresponding operator by $\boldsymbol{H}$. Let $\Psi_{0}\in\mathcal{H}$ be the state at time $t=0$ of a certain quantum system. Then at time $t$ the system is represented by the vector $\Psi\left(t\right)=\boldsymbol{U}_{t}\Psi_{0}$, where $\boldsymbol{U}_{t}=e^{-it\boldsymbol{H}}$, $t\geq 0$, is a unitary operator called the evolution operator. The vector function $\Psi\left(t\right)$ is differentiable if $\Psi\left(t\right)$ is contained in the domain $Dom(\boldsymbol{H})$ of $\boldsymbol{H}$, which happens if at $t=0$, $\Psi_{0}\in Dom(\boldsymbol{H})$, and in this case the time evolution of $\Psi\left(t\right)$ is controlled by the Schrödinger equation $i\frac{\partial}{\partial t}\Psi\left(t\right)=\boldsymbol{H}\Psi\left(t\right)$, where $i=\sqrt{-1}$ and the Planck constant is assumed to be one. For an in-depth discussion of QM, the reader may consult [36]-[37]. In $p$-adic QM, $\mathcal{H=}L^{2}\left(U\right)$, where $U$ is an open compact subset of $\mathbb{Q}_{p}$. For the sake of simplicity, we use $\mathcal{H=}L^{2}\left(\mathbb{Z}_{p}\right)$, where $\mathbb{Z}_{p}=\left\{x\in\mathbb{Q}_{p};\left|x\right|_{p}\leq 1\right\}$ is the unit ball of $\mathbb{Q}_{p}$, which is also a  locally compact additive group.
@@ 3.1. Initial value problems for $p$-adic Schrödinger equations on the unit ball
(3.1) $\left\{\begin{array}[c]{l}\Psi\left(\cdot,t\right)\in L^{2}(\mathbb{Z}_{p})\text{, }t>0;\text{ }\Psi\left(x,\cdot\right)\in C^{1}(0,\infty)\text{, }x\in\mathbb{Z}_{p};\\ \\ i\frac{\partial}{\partial t}\Psi\left(x,t\right)=\boldsymbol{H}\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t>0;\\ \\ \Psi\left(x,0\right)=\Psi_{0}\left(x\right)\in L^{2}(\mathbb{Z}_{p})\text{;}\\ \\ \left\|\Psi\left(\cdot,t\right)\right\|_{2}=\left\|\Psi_{0}\right\|_{2}=1\text{, }t\geq 0.\end{array}\right$.
We assume the existence of a unique solution for (3.1). We denote the unitary group attached to $\boldsymbol{H}$ as $e^{-i\boldsymbol{H}t}$, $t\in\mathbb{R}$.
@@ 3.2. An orthonormal basis for $L^{2}(\mathbb{Z}_{p})$
(3.4) $\Psi\left(x,t\right)=A_{0}\Omega\left(\left|x\right|_{p}\right)e^{-iE_{\text{gnd}}t}+{\displaystyle\sum\nolimits_{r,b,k}}A_{rbk}\Psi_{rbk}\left(x\right)e^{-iE_{rbk}t},$
@@ 3.3. Construction of CTQMCs I
(3.7) $\mathcal{H}_{J}:=\mathit{Span}\left\{e_{j};j\in\mathbb{J}\right\}\text{ is a Hilbert subspace of }L^{2}(\mathbb{Z}_{p}),$
@@
(3.8) $e^{-i\boldsymbol{H}t}\mathcal{H}_{\mathbb{J}}\subset\mathcal{H}_{\mathbb{J}}\text{.}$
@@
(3.9) $\Psi_{v}\left(x,t\right):=e^{-i\boldsymbol{H}t}e_{v}(x)\in L^{2}\left(\mathbb{Z}_{p}\right),$
@@
(3.10) $\pi_{r,v}\left(t\right)=\left|\left\langle e_{r},e^{-i\boldsymbol{H}t}e_{v}\right\rangle\right|^{2}=\left|\left\langle e_{r},\Psi_{v}\right\rangle\right|^{2},$
@@
The presented construction is entirely different from the one given in [2]-[5], since here we use an infinite-dimensional Hilbert space $L^{2}(\mathbb{Z}_{p})$, and we do not use graphs. In contrast, in the references mentioned above, the constructions use finite-dimensional Hilbert spaces and adjacency matrices. On the other hand, we propose to interpret  the state $\infty$ in the previous theorem as an interaction with an external system. We do not present numerical simulations for this type of network here. The main difficulty is the computational implementation of the condition $e^{-i\boldsymbol{H}t}\mathcal{H}_{\mathbb{J}}\nsubseteqq\mathcal{H}_{\mathbb{J}}$.
@@ 4.1. A class of $p$-adic heat equations
Then, $\widehat{\text{ }u}\left(\xi,t\right)=e^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\widehat{u}_{0}\left(\xi\right)$, and setting  $Z_{0}(x,t)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(e^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\right)$, and taking formally the inverse Fourier transform,
$u\left(x,t\right)$ $=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(e^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\widehat{u}_{0}\left(\xi\right)\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(e^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\right)\ast u_{0}\left(x\right)$
@@ 5.1. A class of $p$-adic Schrödinger equations
By performing a Wick rotation $t\rightarrow it$ in (4.1), and taking $\Psi\left(x,t\right)=u\left(x,it\right)$, we obtain the following Schrödinger equations in the unit ball:
(5.1) $\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
@@
(5.2) $\Psi\left(x,t\right)=Z_{0}(x,it)\ast\psi_{0}\left(x\right)\text{ in }\mathcal{D}^{\prime}(\mathbb{Q}_{p})\text{,}$
@@ 6. Discretizations of $p$-adic evolutionary differential equations
We identify $\mathcal{D}_{l}(\mathbb{Z}_{p})$ with the Hilbert space $\mathbb{C}^{p^{l}}$. This identification maps $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\}_{r\in G_{l}}$ into the standard orthonormal basis $\left\{e_{r}\right\}$ $_{r\in G_{l}}$ of $\mathbb{C}^{p^{l}}$. We show that the transition probabilities of the CTQMC, $\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}$, given in (5.3) satisfy $\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|e^{it\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle_{l}\right|^{2}$, for $J,I\in G_{l}$. We use this formula in the numerical simulations given in the next section.

### v2

@@ 1. Introduction
Quantum computing is a computing paradigm that uses quantum physics effects like superposition, interference, and entanglement to achieve computational power beyond classical computing; see, e.g., the recent survey [1], and the references therein. The continuous-time quantum walks (CTQWs) are the quantum counterparts of the classical random walks. CTQWs are a fundamental tool for developing efficient quantum algorithms, simulating complex quantum systems, and potentially achieving quantum advantages in various computational tasks. CTQWs offer potential advantages over classical algorithms, particularly in graph-based problems like spatial search, [2]-[9]. Currently, there is a large variety of quantum physical systems that can be used for implementing quantum walks and their applications. For instance, trapped-ion and trapped-atom systems, bulk-optics systems, and integrated photonics systems. These considerable advances in physical implementations of quantum walks are allowing the implementation of specialized quantum walk computing systems for applications of practical interest, [1], [10].
Ultrametricity in physics means the emergence of ultrametric spaces in physical models. A metric space $(M,d)$ is ultrametric if the distance $d$ satisfies $d(A,B)\leq\max\left\{d(A,C),d(C,B)\right\}$ for any three points $A$, $B,$ $C$ in $M$. Ultrametricity was discovered in the 80s by Parisi and others in the theory of spin glasses and by Frauenfelder and others in physics of proteins; see e.g. [11], [12], see also [13, Chapter 4], and the references therein. In both cases, the space of states of a complex system has a hierarchical structure that plays a central role in the physical behavior of the system.
@@
In [26], the first author showed that a large class of $p$-adic Schrödinger equations is the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). Practically, a discretization of such an equation gives a CTQMC. As a practical result, we construct new types of continuous-time quantum walks (CTQWs) on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. The final goal of the mentioned work was to contribute in the understanding of the role of the hypothesis of the discreteness of space in the foundations of quantum mechanics. The connection between $p$-adic QM and CTQWs show that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory because the Hamiltonians used are non-local operators, and consequently, spooky actions at a distance are allowed. But, this theory is not a mathematical toy. The paradigm asserting that the universe in not locally real, implies that $p$-adic QM allows realism.
@@
$\left\{\begin{array}[c]{l}\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right),\end{array}\right$.
where $\mathrm{i}=\sqrt{-1}$. The solution of is given by $\Psi\left(x,t\right)=Z_{0}\left(x,it\right)\ast\psi_{0}\left(x\right)$, cf. Theorem 11.1, in Appendix C. Now, we construct a quantum Markov chain in $G_{l}$ using the following transition probabilities:
@@
Set $\boldsymbol{J}\varphi=J\ast\varphi-\varphi$, then the operators $\boldsymbol{J}$, $\mathrm{e}^{t\boldsymbol{J}}$, $\mathrm{e}^{it\boldsymbol{J}}$ are limits of finite dimensional operators $\boldsymbol{J}^{\left(l\right)}$, $\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}$, $\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}$, and consequently in the computation of $p_{r,v}(t)$, respectively of $\pi_{r,v}\left(t\right)$, we can use $Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)$, respectively $Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)$, and thus the construction of the above-mentioned Markov chains depends on a matrix $\boldsymbol{J}^{\left(l\right)}$. By suitably selecting this matrix, we construct the CTQWs based on adjacency matrices of graphs, see Appendix E.
@@ 2.3. $p$-Adic QM
In the Dirac-Von Neumann formulation of QM, to every isolated quantum system there is associated a separable complex Hilbert space $\mathcal{H}$ called the space of states. The Hilbert space of a composite system is the Hilbert space tensor product of the state spaces associated with the component systems. Non-zero vectors describe the states of a quantum system from $\mathcal{H}$. Each observable corresponds to a unique linear self-adjoint operator in $\mathcal{H}$. The most important observable of a quantum system is its energy. We denote the corresponding operator by $\boldsymbol{H}$. Let $\Psi_{0}\in\mathcal{H}$ be the state at time $t=0$ of a certain quantum system. Then at time $t$ the system is represented by the vector $\Psi\left(t\right)=\boldsymbol{U}_{t}\Psi_{0}$, where $\boldsymbol{U}_{t}=\mathrm{e}^{-it\boldsymbol{H}}$, $t\geq 0$, is a unitary operator called the evolution operator. The vector function $\Psi\left(t\right)$ is differentiable if $\Psi\left(t\right)$ is contained in the domain $Dom(\boldsymbol{H})$ of $\boldsymbol{H}$, which happens if at $t=0$, $\Psi_{0}\in Dom(\boldsymbol{H})$, and in this case the time evolution of $\Psi\left(t\right)$ is controlled by the Schrödinger equation $\mathrm{i}\frac{\partial}{\partial t}\Psi\left(t\right)=\boldsymbol{H}\Psi\left(t\right)$, where $\mathrm{i}=\sqrt{-1}$ and the Planck constant is assumed to be one. For an in-depth discussion of QM, the reader may consult [36]-[37]. In $p$-adic QM, $\mathcal{H=}L^{2}\left(U\right)$, where $U$ is an open compact subset of $\mathbb{Q}_{p}$. For the sake of simplicity, we use $\mathcal{H=}L^{2}\left(\mathbb{Z}_{p}\right)$, where $\mathbb{Z}_{p}=\left\{x\in\mathbb{Q}_{p};\left|x\right|_{p}\leq 1\right\}$ is the unit ball of $\mathbb{Q}_{p}$; $\mathbb{Z}_{p}$ is a compact additive group.
@@ 3.1. Initial value problems for $p$-adic Schrödinger equations on the unit ball
(3.1) $\left\{\begin{array}[c]{l}\Psi\left(\cdot,t\right)\in L^{2}(\mathbb{Z}_{p})\text{, }t>0;\text{ }\Psi\left(x,\cdot\right)\in C^{1}(0,\infty)\text{, }x\in\mathbb{Z}_{p};\\ \\ \mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=\boldsymbol{H}\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t>0;\\ \\ \Psi\left(x,0\right)=\Psi_{0}\left(x\right)\in L^{2}(\mathbb{Z}_{p})\text{;}\\ \\ \left\|\Psi\left(\cdot,t\right)\right\|_{2}=\left\|\Psi_{0}\right\|_{2}=1\text{, }t\geq 0.\end{array}\right$.
We assume the existence of a unique solution for (3.1). We denote the unitary group attached to $\boldsymbol{H}$ as $\mathrm{e}^{-i\boldsymbol{H}t}$, $t\in\mathbb{R}$.
@@ 3.2. An orthonormal basis for $L^{2}(\mathbb{Z}_{p})$
(3.4) $\Psi\left(x,t\right)=A_{0}\Omega\left(\left|x\right|_{p}\right)\mathrm{e}^{-iE_{\text{gnd}}t}+{\displaystyle\sum\nolimits_{r,b,k}}A_{rbk}\Psi_{rbk}\left(x\right)\mathrm{e}^{-iE_{rbk}t},$
@@ 3.3. Construction of CTQMCs I
(3.7) $\mathcal{H}_{\mathbb{J}}:=\mathit{Span}\left\{e_{j};j\in\mathbb{J}\right\}\text{ is a Hilbert subspace of }L^{2}(\mathbb{Z}_{p}),$
@@
(3.8) $\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}\mathcal{H}_{\mathbb{J}}\subset\mathcal{H}_{\mathbb{J}}\text{.}$
@@
(3.9) $\Psi_{v}\left(x,t\right):=\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}e_{v}(x)\in L^{2}\left(\mathbb{Z}_{p}\right),$
@@
(3.10) $\pi_{r,v}\left(t\right)=\left|\left\langle e_{r},\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}e_{v}\right\rangle\right|^{2}=\left|\left\langle e_{r},\Psi_{v}\right\rangle\right|^{2},$
@@
The presented construction is entirely different from the one given in [2]-[5], since here we use an infinite-dimensional Hilbert space $L^{2}(\mathbb{Z}_{p})$, and we do not use graphs. In contrast, in the references mentioned above, the constructions use finite-dimensional Hilbert spaces and adjacency matrices. On the other hand, we propose to interpret the state $\infty$ in the previous theorem as an interaction with an external system. We do not present numerical simulations for this type of network here. The main difficulty is the computational implementation of the condition $\mathrm{e}^{-\mathrm{i}\boldsymbol{H}t}\mathcal{H}_{\mathbb{J}}\nsubseteqq\mathcal{H}_{\mathbb{J}}$.
@@ 4.1. A class of $p$-adic heat equations
Then, $\widehat{\text{ }u}\left(\xi,t\right)=\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\widehat{u}_{0}\left(\xi\right)$, and setting $Z_{0}(x,t)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\right)$, and taking formally the inverse Fourier transform,
$u\left(x,t\right)$ $=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\widehat{u}_{0}\left(\xi\right)\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}\left(\mathrm{e}^{t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\right)\ast u_{0}\left(x\right)$
@@ 5.1. A class of $p$-adic Schrödinger equations
By performing a Wick rotation $t\rightarrow\mathrm{i}t$ in (4.1), and taking $\Psi\left(x,t\right)=u\left(x,it\right)$, we obtain the following Schrödinger equations in the unit ball:
(5.1) $\left\{\begin{array}[c]{l}\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
@@
(5.2) $\Psi\left(x,t\right)=Z_{0}(x,\mathrm{i}t)\ast\psi_{0}\left(x\right)\text{ in }\mathcal{D}^{\prime}(\mathbb{Q}_{p})\text{,}$
@@ 6. Discretizations of $p$-adic evolutionary differential equations
We identify $\mathcal{D}_{l}(\mathbb{Z}_{p})$ with the Hilbert space $\mathbb{C}^{p^{l}}$. This identification maps
$\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)\right\}_{r\in G_{l}}$
into the standard orthonormal basis $\left\{e_{r}\right\}$ $_{r\in G_{l}}$ of $\mathbb{C}^{p^{l}}$. We show that the transition probabilities of the CTQMC, $\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}$, given in (5.3) satisfy $\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{it\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle_{l}\right|^{2}$, for $J,I\in G_{l}$. We use this formula in the numerical simulations given in the next section.